can someone please solve this one...
use Gauss-Jordan elimination and solve this
3x - y = 5
4x + 2 = 10

Answer by venugopalramana(3286)

can someone please solve this one...
use Gauss-Jordan elimination and solve this
3x - y = 5
4x + 2 = 10
AUGMENTED MATRIX IS
3......-1.......5
4.......2.......10
NR1=R1/3
1......-1/3......5/3
4........2........10
NR2=R2-4*R1
1.......-1/3.............5/3
0.......2+4/3=10/3.......10-20/3=10/3
NR2=R2*3/10
1.......-1/3..............5/3
0.........1................1
NR1=R1+R2/3
1.............0............5/3+1/3=2
0.............1...............1
HENCE X=2
Y=1